the greater of two numbers is one more than twice the smaller. three times the greater exceeds 5 times the smaller by 10. Find the numbers.
Found 2 solutions by edjones, checkley77:
Answer by edjones(8007)

Let x=greater, y=lesser
x=2y+1
3x=5y+10
.
3(2y+1)=5y+10
6y+3=5y+10
y=7
x=15
.
Ed

Answer by checkley77(12844)

X=2Y+1
3X=5Y+10
3(2Y+1)=5Y+10
6Y+3=5Y+10
6Y-5Y=10-3
Y=7 ANS FOR THE SMALLER NUMBER.
X=2*7+1=14+1=15 ANS. FOR THE LARGER NUMBER.
PROOF:
3*15=5*7+10
45=35+10
45=45